OpenStudy (anonymous):

When you solve an equation, how do you know when to add and when to subtract? You may include an example if you need to

OpenStudy (thatgirl3390):

4x-2=0 You add 2 to both sides because it's negative 4x=2

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

you add or subtract considering the possibility that either of variables is eliminated suppose x + y =3 and 2x-y =0 here if you add you get 3x =3 n x=1 but if you subtract you get 2y-x= 3 depends on how u want it to be solved